# Supersymmetric scale-separated AdS3 vacua of type IIB

### v1

@@ 5 Summary
In this paper, supersymmetric AdS3 vacua of type IIB string theory with the parametric scale separation feature are constructed. These arise as compactifications on seven-dimensional solvmanifolds with co-closed $G_{2}$-structure, which can be seen as twisted seven-tori. The orbifolds and orientifold are carefully chosen such that all one- and two-forms are projected out and that, in one case, the geometry admits four sets of O5-planes and in another case only one. This means that some components of the $F_{3}$ Bianchi identity are not restricted and lead to unbounded fluxes. In principle, this could also be achieved by having seven sets of O5-planes as in ref. Emelin:2021gzx (so orbifolds without shifts), but it requires including D5-branes to cancel the O5-charge in some directions. It also leads to orbifold singularities for which it is unknown how to resolve them, even for the flat seven-torus. Additionally, such orbifolds also generate an O9-plane image and therefore the model resides in type I string theory instead (after adding 32 D9-branes). Working with the different orbifolds of this paper does not require the inclusion of D5-branes and keeps the setup in type IIB. Therefore, the solutions presented here are, as far as the author knows, the first supersymmetric, and parametrically scale-separated solutions of type IIB string theory at large volumes of the cycles and weak coupling, in the limit where the orientifold planes are smeared. Furthermore, the solution with only one set of O5-planes can be understood as a supersymmetric cousin of one of the vacua of ref. Arboleya:2024vnp through the skew-whiffing procedure. Moreover, the mass spectra of all the supersymmetric solutions, at tree level, lead to operators in the putative dual field theory with integer conformal dimensions, very similar to the investigations of refs. Conlon:2020wmc ; Apers:2022vfp ; Ning:2022zqx for 4d scale-separated vacua and contrary to the 3d scale-separated type IIA vacua, as explored in ref. Apers:2022zjx . The 3d solutions presented here share even more similarities with their 4d type IIA counterparts, especially the ones in massless type IIA: they rely on intersecting O-planes to stabilise the moduli and the internal manifold has non-trivial intrinsic torsion. The fact that the O-planes seem to intersect might be a worry as such setups are not well understood, also not at the world-sheet level. However, ref. Junghans:2023yue showed that for the 4d DGKT vacua DeWolfe:2005uu ; Camara:2005dc on the specific $\mathbb{T}^{6}/\mathbb{Z}_{3}^{2}$ orbifold, the four sets of O6-planes are not intersecting after blowing up the orbifold singularities. Although it is not entirely clear how the orbifold singularities of the twisted tori should be resolved, one could expect that a blow-up procedure, similarly for flat $\mathbb{T}^{7}$’s, might do so. The blow-up moduli could then be stabilised by balancing local RR three-form flux against the bulk seven-form flux, analogous to the blow-up procedure in DGKT DeWolfe:2005uu . A second reason for worry might be that the O-planes are smeared over the directions transverse to them, not being treated as the local sources that they are. Although computing the local O-plane backreaction for the 3d vacua is a computationally contrived problem here, previous successful attempts for other AdS vacua seem to indicate that there are no fundamental objections to it in principle. Indeed, in massive type IIA, refs. Baines:2020dmu ; Junghans:2020acz ; Marchesano:2020qvg ; Emelin:2022cac ; Andriot:2023fss ; Emelin:2024vug showed that on various backgrounds without intrinsic torsion, the O-plane backreaction is small and can be computed perturbatively, at first order. Additionally, refs. Cribiori:2021djm ; VanHemelryck:2024bas argued the same for backgrounds of massless type IIA with intrinsic torsion, where calculating the backreaction was essential to lift the solutions to M-theory. Finally, although the solutions in this paper rely on a specific class of seven-dimensional solvmanifolds, it would be interesting to explore other co-closed $G_{2}$-structures that yield similar scale-separated solutions in type IIB string theory. Indeed, the ideas presented here suggest that the search for new scale-separated vacua is not over yet.
@@ Acknowledgements
I want to thank Adolfo Guarino, Matteo Morittu, Thomas Van Riet, Max Hübner, Magdalena Larfors and especially Álvaro Arboleya and George Tringas for interesting discussions. I am financially supported by the Olle Engkvists Stifelese.

### v3

@@ 4 Summary
This paper investigates compactifications of type IIB string theory on seven-dimensional manifolds equipped with co-closed $G_{2}$-structures that arise as $\mathbb{Z}_{2}^{3}$ orbifolds of twisted tori and preserve minimal supersymmetry. More specifically, the geometries considered here are one solvmanifold and all nilmanifolds compatible with the orbifold. All the solutions require orientifolds, such that the tadpole from the $F_{3}$ Bianchi identity can be cancelled by O5-planes wrapping the right three-cycles. Since the amount of orientifold planes cannot be tuned in string theory, some components of the $F_{3}$ flux must be fixed by the tadpole, whereas others are unbounded. It is by sending the latter fluxes and the seven-form flux to large values that the well-controlled and scale-separated regime can be achieved for these vacua. Crucially, these backgrounds do not require O5-planes to be wrapped along all seven three-cycles. Shifts were implemented in the $\mathbb{Z}_{2}^{3}$ orbifold actions such that only the necessary intersecting orientifold images are generated. Note that in principle, the $F_{3}$ Bianchi identity can also be solved by wrapping O5-planes along all seven three-cycles as in ref. Emelin:2021gzx (realised by orbifolds without shifts), but by also including D5-branes to cancel the O5-charge along some cycles. It also leads to orbifold singularities for which it is unknown how to resolve them, even for the ordinary seven-torus. Additionally, such orbifolds also generate an O9-plane image and therefore the model resides in type I string theory instead (after adding 32 D9-branes). Working with the different orbifolds of this paper does not require the inclusion of D5-branes and keeps the setup in type IIB string theory.
On the solvmanifolds, the large flux limit leads to solutions in the well-controlled regime, i.e. with weak string coupling and large cycles in string units, and with a hierarchy between the AdS scale and the cycle volumes. However, these cycle volumes do not seem to provide good proxies for the KK scale, which is usually estimated by the lowest eigenvalue of the scalar Laplacian. Therefore it cannot be concluded that these solutions are scale-separated. The reason is that preliminary analysis shows that the lowest eigenvalue is given by the biggest length scale of the manifold ($L_{7}$) which, through supersymmetry, behaves parametrically the same as the AdS radius. Note that the same issue arises in the setup of refs. Arboleya:2024vnp ; Arboleya:2025ocb although they are non-supersymmetric. Indeed, the solution with only one set of O5-planes can be understood as a supersymmetric cousin of one of the vacua of ref. Arboleya:2024vnp ; Arboleya:2025ocb through the skew-whiffing procedure. The aforementioned issue does not arise for nilmanifolds, as the AdS- and internal curvature scale decouple from the KK scale Andriot:2016rdd ; Andriot:2018tmb . In this paper, only for the nilmanifolds with names $\mathfrak{n}_{2}$, $\mathfrak{n}_{4}$, $\mathfrak{n}_{6}$, 37B1, 37C and 37D1, were solutions found where there exists a (partially) homogeneous large flux-limit which leads to large radii, small string coupling and an AdS radius that becomes parametrically larger than any of the radii, and consequentially the lowest eigenvalue of the Laplacian, as the results of refs. Andriot:2016rdd ; Andriot:2018tmb indicate.
Therefore, the solutions presented here are, as far as the author knows, the first supersymmetric and parametrically scale-separated solutions of type IIB string theory at large cycle volumes and weak coupling, in the limit where the orientifold planes are smeared.
Interestingly, the mass spectra of some of the solutions, that is, on the solvmanifolds and nilmanifold $\mathfrak{n}_{2}$, lead to operators in the putative dual field theory with integer conformal dimensions at tree level. The supersymmetric solutions are fully homogeneous under a rescaling of the fluxes. As a consequence, the masses (in AdS units) do not depend on those fluxes. Moreover, all fluxes that are sent to infinity are ‘trivially unbounded’, i.e. they are all independent. This is very similar to the investigations of refs. Conlon:2020wmc ; Apers:2022vfp ; Ning:2022zqx for 4d scale-separated vacua. However, on all the other backgrounds, the conformal dimensions are not integers. An interesting observation is that for all those vacua, there are scalar fields that have masses in AdS units that are dependent on the fluxes. For all of them except for $\mathfrak{n}_{4}$, there are also ‘non-trivially unbounded fluxes’, meaning that although they can be sent to infinity, they have to satisfy a constraint. For these cases, there exist flux limits for which some of the scalars become very light compared to the AdS scale. For the 3d massive type IIA vacua with scale separation that also features non-trivially unbounded fluxes, the conformal dimensions were also found not to be all integers, as noted in ref. Apers:2022zjx , and it would be interesting to investigate whether there also exist flux limits where some scalars become light. Moreover, variants on this model with trivially unbounded fluxes Farakos:2025bwf , albeit having massless moduli, also have integer conformal dimensions. Therefore, they seem to support the idea that there is a relation between setups with trivially unbounded fluxes and the presence of integer conformal dimensions and vice versa. Moreover, it would be interesting to explore further holographic aspects of these solutions in more detail. A starting point could be provided by the recently proposed flux backtracking algorithm Apers:2025pon . Especially the solution on the nilmanifold $\mathfrak{n}_{2}$ seems promising, as this is a compactification to $\mathrm{AdS}_{3}\times\mathrm{Nil}_{3}\times\mathbb{T}^{4}$ with $F_{3}$ and $F_{7}$ flux and (intersecting) O-planes, which is potentially dual to the world-volume theory of a non-trivial D1-D5 system. Furthermore, this construction also opens up the possibility of going beyond the twisted torus Ansatz by considering compactifications of the type $\mathrm{AdS}_{3}\times\mathrm{Nil}_{3}\times\mathcal{M}_{4}$, where $\mathcal{M}_{4}$ is a four-dimensional Ricci-flat manifold. This is left for future research.
@@
The 3d solutions presented here share some similarities with their 4d type IIA counterparts, especially the ones in massless type IIA: they rely on intersecting O-planes to stabilise the moduli, and the internal manifold has non-trivial intrinsic torsion. The fact that the O-planes seem to intersect might be a worry, as such setups are not well understood, also not at the world-sheet level. However, ref. Junghans:2023yue showed that for the 4d DGKT vacua DeWolfe:2005uu ; Camara:2005dc on the specific $\mathbb{T}^{6}/\mathbb{Z}_{3}^{2}$ orbifold, the four sets of O6-planes are not intersecting after blowing up the orbifold singularities. Although it is not entirely clear how the orbifold singularities of the twisted tori should be resolved, one could expect that a blow-up procedure, similar to ordinary $\mathbb{T}^{7}$’s, might do so. The blow-up moduli could then be stabilised by balancing local RR three-form flux against the bulk seven-form flux, analogous to the blow-up procedure in DGKT DeWolfe:2005uu . A second reason for worry might be that the O-planes are smeared over the directions transverse to them, not being treated as the local sources that they are. Although computing the local O-plane backreaction for the 3d vacua is a computationally contrived problem here, previous successful attempts for other AdS vacua seem to indicate that there are no fundamental objections to it in principle. Indeed, in massive type IIA, refs. Baines:2020dmu ; Junghans:2020acz ; Marchesano:2020qvg ; Emelin:2022cac ; Andriot:2023fss ; Emelin:2024vug showed that on various backgrounds without intrinsic torsion, the O-plane backreaction is small and can be computed perturbatively, at first order. Additionally, refs. Cribiori:2021djm ; VanHemelryck:2024bas argued the same for backgrounds of massless type IIA with intrinsic torsion, where calculating the backreaction was essential to lift the solutions to M-theory. Altogether, these results provide evidence that parametrically scale-separated, supersymmetric AdS3 vacua can be constructed in type IIB string theory at large volume and weak coupling in the presence of intersecting orientifolds. Although all these solutions relied on specific $\mathbb{Z}_{2}^{3}$ orbifolds of seven-dimensional nilmanifolds, it would be interesting to explore other options that yield similar scale-separated solutions in type IIB string theory, like different orbifolds or geometries. Indeed, the ideas presented here suggest that the search for new scale-separated vacua is not over yet.
@@ Acknowledgements
I want to thank Adolfo Guarino, Matteo Morittu, Thomas Van Riet, Max Hübner, Magdalena Larfors and especially Álvaro Arboleya and George Tringas for interesting discussions. I am financially supported by the Olle Engkvists Stiftelse.
Appendix A Laplacian spectrum of the ISO(2) solvmanifold
This appendix briefly discusses the eigenfunctions and eigenvalues of the scalar Laplacian of the 3d ISO(2) solvmanifold. [^9] The full spectrum on the related 3d ISO(1,1) solvmanifold has been studied in ref. MR2217282 . These can easily be generalised to the 7d solvmanifold used in this paper. For this solvmanifold, a twisted torus metric as in eq. (2) can be used with the co-frame one-forms as in eq. (33):
$\begin{split}\mathrm{d}s_{3}^{2}=&L_{1}^{2}\left(\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}-\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\right)^{2}+L_{2}^{2}\left(\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}+\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\right)^{2}\\ &L_{7}^{2}\left(\mathrm{d}x^{7}\right)^{2}\,.\end{split}$ (138)
With this metric, one can take the following Ansatz for the eigenfunctions:
$Y_{klm}=\mathrm{e}^{2\pi i(kx^{1}+lx^{2})}Z_{m}(x^{7})\,.$ (139)
Plugging this into the eigenvalue equation
$(\Delta-\lambda_{klm})Y_{klm}=0\,$, (140)
and constructing the Laplacian with the metric (138), one finds that the eigenvalue equation reduces to the Mathieu equation for $Z_{m}$ NIST:DLMF :
$Z_{m}^{\prime\prime}+\left[a_{klm}-2q_{kl}\cos\left(2(\omega_{a}x^{7}+\varphi_{kl})\right)\right]Z_{m}=0\,$, (141)
where the parameters $a_{klm}$, $q_{kl}$ and $\varphi_{kl}$ are given by:
$a_{klm}=-\frac{L_{7}^{2}}{\omega_{a}^{2}}\left(\frac{(2\pi)^{2}(k^{2}+l^{2})(L_{1}^{2}+L_{2}^{2})}{2L_{1}^{2}L_{2}^{2}}+\lambda_{klm}\right),\quad q_{kl}=-\frac{L_{7}^{2}}{\omega_{a}^{2}}\left(\frac{(2\pi)^{2}(k^{2}+l^{2})(L_{1}^{2}-L_{2}^{2})}{4L_{1}^{2}L_{2}^{2}}\right)$
$\varphi_{kl}=\frac{1}{2}\arctan\left(\frac{2kl}{k^{2}-l^{2}}\right)\,.$ (142)
The solutions to this equation are periodic in $x^{7}$ only for specific values of $a_{klm}$, which are called the Mathieu characteristic values, also named $a_{m}(q)$ and $b_{m}(q)$, such that
$a_{klm}=a_{m}(q_{kl})\quad\text{or}\quad b_{m}(q_{kl})\,$, (143)
where $m$ is an integer. This discretises the eigenvalue spectrum. The eigenfunctions are then written as (co)sine elliptic functions as follows NIST:DLMF :
$Z_{m}=\{\mathrm{ce}_{m}(\omega_{a}x^{7}+\varphi_{kl},q_{kl}),\mathrm{se}_{m}(\omega_{a}x^{7}+\varphi_{kl},q_{kl})\}\,$, (144)
where $\mathrm{ce}_{m}$ and $\mathrm{se}_{m}$ are even and odd functions respectively. The eigenvalues of $\mathrm{ce}_{m}(x,q)$ and $\mathrm{se}_{m}(x,q)$ are given by $a_{m}(q)$ and $b_{m}(q)$ respectively, with the caveat that $\mathrm{se}_{0}(x,q)$ and $b_{0}(q)$ do not exist. Additionally, when $m$ is even, then the functions have a periodicity of $\pi$, whereas when they are odd they have antiperiodicity $\pi$. However, the case of interest is when $k=l=0$, hence when the eigenfunctions only depend on $x^{7}$. Then, the eigenvalue equation (141) reduces to the simple harmonic oscillator and the eigenvalues are:
$\lambda_{00m}=-\frac{m^{2}\omega_{a}^{2}}{L_{7}^{2}}\,.$ (145)
Upon ignoring that $\varphi_{00}$ is ill-defined, this indeed agrees with the Mathieu eigenfunctions when $k=l=0$, as it implies $q_{kl}=0$ (which also happens when $L_{1}=L_{2}$). In that case, $a_{m}$ and $b_{m}$ become analytic and simple:
$a_{m}(0)=b_{m}(0)=m^{2},$ (146)
and the (co)sine elliptic functions become ordinary (co)sine functions, as expected. So the tower of eigenvalues given by eq. (145) only depends on $L_{7}$. Because $L_{7}$ is the largest radius, it sets the lowest eigenvalue and the KK scale although $L_{7}$ decouples from the cycle volumes. Ordinary $\mathbb{Z}_{2}^{n}$ orbifolds cannot project out this tower of eigenfunctions. Alternatively, when $k\neq 0$ or $l\neq 0$ and $L_{2}\gtrsim L_{1}$, then $q_{kl}\neq 0$, and because $L_{7}$ is assumed to be much larger than $L_{1}$ and $L_{2}$, $q_{kl}$ is large (as $L_{2}-L_{1}=O(1)$). In this limit, both $a_{mkl}$ and $b_{mkl}$ take the same form at leading order for all $m$:
$a_{mkl}=-2q_{kl}+2(2m+1)\;\sqrt{q_{kl}}-\frac{2m(m+1)+1}{4}+O\left(q_{kl}^{-1/2}\right)\,$, (147)
This implies that the eigenvalues at leading order are
$\lambda_{klm}$ $=-\frac{\omega_{a}^{2}}{L_{7}^{2}}a_{klm}-\frac{(2\pi)^{2}(k^{2}+l^{2})(L_{1}^{2}+L_{2}^{2})}{2L_{1}^{2}L_{2}^{2}}$
$\approx 2\frac{\omega_{a}^{2}}{L_{7}^{2}}q_{kl}-\frac{(2\pi)^{2}(k^{2}+l^{2})(L_{1}^{2}+L_{2}^{2})}{2L_{1}^{2}L_{2}^{2}}=-\frac{(2\pi)^{2}(k^{2}+l^{2})}{L_{2}^{2}}\,$, (148)
which, up to corrections, does not depend on $L_{7}$, and are bigger in size than $\lambda_{00m}$ as $L_{7}\gg L_{2}$. Finally, it is clear that all these aforementioned eigenfunctions and eigenvalues are also included in the spectrum of the Laplacian of 7d solvmanifold used in this paper. Indeed, the differential equation (141) would involve a sum of cosines over the 3 different ISO(3) copies, but one can look at modes involving only one copy, as what is done here.
Appendix B Mass spectra of the nilmanifold compactifications
In this appendix, the mass spectra of the different nilmanifold compactifications of section 3 are discussed.
Nilmanifold $\mathfrak{n}_{2}$
This nilmanifold is the simplest of the list, and the scalar masses do not depend on the fluxes as the setup is fully homogeneous. The scalar masses are computed to be
$m^{2}L_{\mathrm{AdS}}^{2}=\{120,8,8,8,8,8,8,8\}\,$, (149)
and this leads to integer conformal dimensions, as discussed in the main text.
Nilmanifold $\mathfrak{n}_{4}$
This nilmanifold is special in the sense that the compactification is only partially homogeneous under flux rescalings, but does not have non-trivially unbounded fluxes. There are 6 out of 8 scalar masses that are flux-independent and are all the same with
$m^{2}L_{\mathrm{AdS}}^{2}=8.$ (150)
The other two scalars have mass that depends on $\rho$, but there were only a finite number of consistent values for $\rho$, 16 to be precise. The masses squared are plotted in Figure 1.
Figure 1: Masses squared for nilmanifold $\mathfrak{n}_{4}$ for 16 different values of $\rho$, left and right for a scalar with negative and positive mass squared, respectively.
Nilmanifold $\mathfrak{n}_{6}$
For this nilmanifold, there are non-trivially unbounded fluxes, and the solutions are parametrised here by a parameter $\tau$ introduced first in eq. (91). Irrespective of $\tau$, the trace of the mass matrix is 80, meaning that the total sum of the masses should be 80. For this nilmanifold, 3 out of 8 scalar masses are all the same with
$m^{2}L_{\mathrm{AdS}}^{2}=8.$ (151)
The other 6 scalars have masses that depend on $\rho$ and $\tau$, which are plotted in Figure 2 corresponding to the two branches of $\rho$ (see eq. (96)).
Figure 2: Masses squared for nilmanifold $\mathfrak{n}_{6}$ dependent on $\tau$, for the - branch (left) and ++ branch (right). For all lines, $a=1$, except for the red and purple, which are normalised with $a=8$ and $a=48$ respectively.
In the large $\tau$ limit, three of the masses asymptote to 0, whereas the other two go to the values 8 and 48.
$m^{2}L_{\mathrm{AdS}}^{2}\underset{\tau\to\infty}{=}\left\{48,8,8,8,8,0,0,0\right\}+O(\tau^{-1})\,.$ (152)
Even at large $\tau$, one can obtain scale separation in the well-controlled regime, as long as the flux condition (98) is replaced by
$f_{3}^{2}\tau^{2}\ll f_{7}\ll f_{3}^{4},\qquad\tau^{2}\ll f_{3}\,.$ (153)
Nilmanifold 37B1
For nilmanifold 37B1, the trace of the mass matrix is also equal to 80. Two out of eight scalar masses are the same and independent of $\tau$ (as introduced in eq. (102):
$m^{2}L_{\mathrm{AdS}}^{2}=8.$ (154)
The other 5 scalars have mass that depends on $\rho$, which are plotted in Figure 3, for both the ++ and - branches of eq. (107).
Figure 3: Masses squared for nilmanifold 37B1 dependent on $\tau$, for the - branch (left) and ++ branch (right). For all lines, $a=1$, except for the brown one, which is normalised with $a=4$. The plot of the - branch starts from $\tau=1$, as the solution is only consistent for this branch when $\tau>1$.
In the large $\tau$-limit, three scalars become light, and the asymptotic values of all scalars are:
$m^{2}L_{\mathrm{AdS}}^{2}\underset{\tau\to\infty}{=}\left\{48,8,8,8,8,0,0,0\right\}+O(\tau^{-1})\,.$ (155)
A well-controlled and scale-separated regime still exists in this limit, as long as the conditions on the fluxes (109) are modified to
$f_{7}\tau^{2}\ll F_{235}^{4}\ll f_{7}^{2}\tau^{-4},\tau^{2}\ll f_{7}\,.$ (156)
Nilmanifold 37C
Many qualitative features of the mass spectra for the previous nilmanifolds persist for nilmanifold 37C: the trace of the mass matrix is equal to 80, and there are also 3 out of 8 scalar masses that are all the same and independent of $\tau$ and $\rho$ as introduced in eq. (113):
$m^{2}L_{\mathrm{AdS}}^{2}=8.$ (157)
The other 5 scalars have masses that depend on $\rho$ and $\tau$, and are plotted in Figure 4 for the two branches of (117).
Figure 4: Masses squared for nilmanifold 37C dependent on $\tau$, for the - branch (left) and ++ branch (right). For all lines, $a=1$, except for the red and purple, which are normalised with $a=8$ and $a=48$ respectively.
In the large $\tau$ limit, three of the masses asymptote to 0, whereas the other two go to the values 8 and 48.
$m^{2}L_{\mathrm{AdS}}^{2}\underset{\tau\to\infty}{=}\left\{48,8,8,8,8,0,0,0\right\}+O(\tau^{-1})\,.$ (158)
Also in this limit, one can obtain scale-separated and well-controlled solutions, as long as the flux limit (119) is modified to
$F_{235}^{2}\tau^{4}\ll f_{7}\ll F_{235}^{4}\tau^{2}\,,\qquad\tau^{6}\ll f_{7}\,.$ (159)
Nilmanifold 37D1
Also, this last nilmanifold shares some features of the mass spectra with the previous nilmanifolds: the trace of the mass matrix is equal to 80, but this time there is only one scalar mass that is independent of $\tau$ and $\rho$ as introduced in eq. (123) with
$m^{2}L_{\mathrm{AdS}}^{2}=8.$ (160)
The other 7 scalars have mass that depends on $\tau$, which are plotted in Figure 5 corresponding to the two branches of (127).
Figure 5: Masses squared dependent on $\tau$, for the - branch (left) and ++ branch (right). For all lines, $a=1$, except for the upper curve in light blue, which is normalised with $a=4$. The green curve (third highest up) has multiplicity 2. For the - branch, $\rho$ is only positive if $\tau>\sqrt{3}$ and hence the plot starts from this value.
In the large $\tau$ limit, three of the masses asymptote to 0, whereas the other four go to the values 8 and 48.
$m^{2}L_{\mathrm{AdS}}^{2}\underset{\tau\to\infty}{=}\left\{48,8,8,8,8,0,0,0\right\}+O(\tau^{-1})\,.$ (161)
Also here, one can obtain scale separation in the well-controlled regime for large $\tau$, as long as the flux condition (129) is replaced by
$F_{235}^{2}\tau^{6}\ll f_{7}\ll F_{235}^{4}\tau^{2}\,,\qquad\tau^{2}\ll F_{235}\,.$ (162)